OpenStudy (anonymous):

limit as x approaches 0 of (cosx-1+.5x^2)/x^4

OpenStudy (jdoe0001):

\(\bf lim_{x\to 0}\ \cfrac{cos(x-1)+0.5x^2}{x^4}?\qquad lim_{x\to 0} \cfrac{cos(x-1)+\frac{1}{2}x^2}{x^4}?\)

OpenStudy (anonymous):

yea thats the same thing, how do i solve it?

OpenStudy (anonymous):

@drafia are you sure it's not \[\frac{\cos x-1+\dfrac{1}{2}x^2}{x^4}~~?\]